Adipic acid, HOOC(CH2)4COOH, is among the top 50 manufactured chemicals in the United States (nearly 1 million metric tons annually). Its chief use is in the manufacture of nylon. It is a diprotic acid having Ka1=3.9×10−5 and Ka2=3.9×10−6. A saturated solution of adipic acid is about 0.10 M HOOC(CH2)4COOH.

A) Calculate the concentration of [H3O+][H3O+] in this solution.

Express your answer to two significant figures and include the appropriate units.

B) Calculate the concentration of [HOOC(CH2)4COO−][HOOC(CH2)4COO−] in this solution.

Express your answer to two significant figures and include the appropriate units.

C) Calculate the concentration of [HOOC(CH2)4COOH][HOOC(CH2)4COOH] in this solution.

Express your answer to two significant figures and include the appropriate units.

D) Calculate the concentration of [−OOC(CH2)4COO−][−OOC(CH2)4COO−] in this solution.

Express your answer to two significant figures and include the appropriate units.

E) Calculate the concentration of [OH−][OH−] in this solution.

Express your answer to two significant figures and include the appropriate units.
Answer #1

та +нт — Соон(cң, 2,Coo Соон (си, ), соон iniil: o-10 lat equil: 0.2o- x - - ka, = 3-935 х2 -10-х о - к? + 3.9ҳlo-x - 0-39x10tu .00 01955 m @ ® © 0 (H120*) = x+y = 0.0019589 on [ cool (Che),600-] = x = 0.001955 m. [ coon (che de COOH) = 0.-32 = 0.098